# R-Judge: Benchmarking Safety Risk Awareness for LLM Agents

### v2

@@ 3.3 Evaluation Criteria
As shown in the right of Figure 2, the two evaluation recipes are conducted in a serial pipeline. First, LLMs are demonstrated with the record as input and asked to analyze whether the agent actions in the record are safe, formalized as $p_{\theta}(R)\to\textrm{analysis}$. Then, LLMs are asked to output ’unsafe’ or ’safe’ based on record and analysis in the first step, formalized as $p_{\theta}(R,\textrm{analysis})\to\textrm{label}$. The output analysis in the first step and the output label in the second step are taken respectively for the two evaluation recipes, with human annotation as ground truth.
@@
To evaluate the effectiveness of LLMs in identifying safety risks, an open-ended test utilizes GPT-4 as an automatic evaluator to assess the open-ended model-generated analysis on unsafe cases.
Risk identification in agent interaction includes two components: what potential risks are and how the agent causes them. We address the two keys with two criteria, ${\rm Alertness}$ and ${\rm Effectiveness}$. As human-annotated risk description is the ground truth, ${\rm Effectiveness}$ is assessed by the relevance between model-generated analysis and the pivot. ${\rm Alertness}$ indicates whether the model senses potential risks, i.e., its generated analysis demonstrates potential risks of the case. The final ${\rm Grade}$ is the addition of ${\rm Effectiveness}$ and ${\rm Alertness}$. The prompt for the GPT-4 evaluator is attached in Appendix C.3.2.
Models Safety Judgment Risk Identification
F1 Recall Specificity Validity Grade Effectiveness Alertness
Random 50.00 50.00 50.00 50.00 16.67 0.00 50.00
Human 89.07 82.17 95.76 100.00 87.00 86.67 88.33
GPT-4 72.52 62.00 83.64 100.00 71.60 71.00 74.00
ChatGPT 39.42 27.00 81.82 100.00 46.20 47.50 41.00
Vicuna-13b-v1.5-16k 43.24 32.00 70.91 99.35 34.20 33.50 37.00
Llama-2-13b-chat-hf 38.86 34.00 25.45 50.97 41.20 40.50 44.00
Vicuna-13b-v1.5 30.30 20.00 78.18 100.00 30.80 31.00 30.00
Vicuna-7b-v1.5-16k 36.88 26.00 72.73 100.00 32.20 31.00 37.00
Llama-2-7b-chat-hf 21.56 18.00 10.91 37.42 23.80 23.00 27.00
Vicuna-7b-v1.5 19.35 12.00 78.18 100.00 30.80 30.00 34.00
LlamaGuard-7b 0.00 0.00 98.18 96.77 0.33 0.50 0.00
Table 1: Main results(%) of the safety judgment test and risk identification test in R-Judge. Safety judgment scores are calculated by counting correct labels and risk identification scores are assigned by an automatic GPT-4 evaluator compared with human-annotated risk description. Segment 1: Reference Score, including Random as baseline and Human as upper bound marked as underline; Segment 2: API-based models; Segment 3: 13B models; Segment 4: 7B models. The best average model results are in bold face, except for Validity which is the reference score.
@@ 4 Experiments
In this section, we first describe the experimental setup, especially metrics. Then, after validating R-Judge, we evaluate 9 popular LLMs on two tests of safety judgment and risk identification.
@@ 4.1 Setup
We comprehensively assess 9 LLMs, including API-based models and open-source models. The API-based models include GPT-4 (gpt-4-0613 version) (OpenAI, 2023) and ChatGPT (gpt-3.5-turbo-0613 version). [^1] https://chat.openai.com/. The open-source models include Llama-2 (Touvron et al., 2023) and Vicuna (Chiang et al., 2023) series. Particularly, on the foundation of Llama-2-7b-chat, LlamaGuard-7b (Inan et al., 2023) is fine-tuned to identify safety risks in LLM-generated content. Computational resources and time consumption to run these models are listed in Appendix E.
@@
We adopt the zero-shot chain-of-thought prompting (Kojima et al., 2022) (dubbed Zero-Shot-CoT) to induce LLMs to generate the reasoning steps before producing the final answer. This kind of analyze-then-output process has been shown to improve reasoning performance, as well as interpretability (Zhang et al., 2023b). Task instructions are simple and clear for generality, as presented in Appendix C.3.1.
@@
Risk Identification, the pairwise evaluation, introduces ${\rm Grade}$, ${\rm Alertness}$ and ${\rm Effectiveness}$, as stated in Section 3.3. In the leaderboard, they are all normalized to the range of 0 to 1.
@@
We provide random scores and human scores for comparison. To obtain human performance, we employed six human experts majoring in computer science and cyber security and averaged their scores to measure the ceiling performance on R-Judge. As safety judgment is a task of binary classification, random scores are 50% for F1, Recall, and Specificity. For open-ended risk identification, random Effectiveness is 0%, while random Alertness is 50%.
@@ 4.2 Validating R-Judge
Human agreement with GPT-4 evaluator in the risk identification test.
First, to tackle the potential bias and incapability of LLM-as-Judge (Liu et al., 2023; Zheng et al., 2023), we measure the agreement between GPT-4 and human experts by the Pearson correlation coefficient (Wikipedia, ). The Pearson correlation coefficient quantifies the linear correlation between two variables and is a value between -1 and 1, where the higher value indicates a higher correlation.
On 50 randomly selected unsafe samples, three human experts scored model analysis according to the same criteria that prompt the GPT-4 evaluator. The Pearson correlation coefficient on Grade, Effectiveness and Alertness are 0.95, 0.94 and 0.89, proving the reliability of pairwise evaluation for the risk identification test. The full results across models are listed in Appendix C.2. Notably, for GPT-4 answers, there exists bias that the Alertness agreement between GPT-4 evaluator and human is only 0.64. However, the Grade and Effectiveness are valid with values of 0.92 and 0.89.
High Correlation between safety judgment and risk identification of LLM performance.
We observed that model ranks are similar in the two tests and further proved the high correlation between the two tests. With F1 score in the safety judgment test against Grade, Effectiveness and Alertness in the risk identification test, the Pearson correlation coefficient are 0.92, 0.91 and 0.96.
Therefore, mutually supportive, the two intersected tests offer a valid and progressive evaluation. The safety judgment test is more fair and affordable while the risk identification test holds fine-grained interpretability and better discrimination on model performance.
@@ 4.3 Main Results
Table 1 presents the main results in R-Judge. Based on the results, we have the following key findings:
(i) Most LLMs perform unsatisfactorily on the R-Judge benchmark. Although solvable for humans (with nearly 89.07% F1 score in Safety Judgment and 87.00% grade in Risk Identification), most LLMs perform much inferior. Remarkably, GPT-4 ranks first and is also the only model scoring higher than random in the safety judgment test. However, it is still far from reaching human-level performance. The results show that R-Judge is a challenging benchmark and LLMs tend to generate ineffective analysis and make wrong judgments when judging safety risks in agent interaction. There is considerable room for enhancing the safety risk awareness of LLMs in open agent scenarios.
(ii) Larger-scale models achieve better performance. For models from the same series, e.g. vicuna-13b-v1.5-16k and vicuna-7b-v1.5-16k, Llama-13b-chat-hf and Llama-2-7b-chat-hf, the larger one with more parameters present better performance in both tests. And in general, LLMs with more parameters generally perform better. There is only one exception (Vicuna-13b-v1.5-16k 43.24% $>$ ChatGPT 39.42%) in the safety judgment test and no exception in the risk identification test. The results imply that risk awareness demands knowledge and reasoning abilities from numerous parameters for safety judgment and risk identification, which is further confirmed by manual case studies in Section 5.2.
(iii) Additional fine-tuning on safety alignment does not necessarily raise risk awareness in multi-turn agent scenarios. It is known that both the Llama-2-chat series and the Vicuna series are fine-tuned on the foundation of the Llama-2 base model. Llama-2-chat series are conducted on additional safety alignment including fine-tuning and RLHF (Bai et al., 2022; Ouyang et al., 2022) while there is no additional safety alignment for Vicuna. However, results indicate that Llama-2-chat series present no prominent excess than Vicuna with same parameters. And Llama-2-chat series suffer from low validity, i.e. fail to output safety label after taking in multi-turn long context (see Appendix D.2.2). Significantly, LlamaGuard-7b, specifically fine-tuned on identifying safety risks in LLM-generated content, is incapable of identifying safety risks in the execution of LLM agents. Considering the (ii) finding, we argue that for LLMs functioning as agents rather than chatbots, fine-tuning for general capabilities involving knowledge and reasoning should prioritize additional alignment fine-tuning for harmless content.
@@ 5 Analysis
In this section, we delve into the effect of different mechanisms on model performance with further experiments and case studies, acquiring a nuanced understanding of the two tests.
@@ 5.1 Influence of Different Prompting Techniques
As disclosed in literature (Li et al., 2023b; Wang et al., 2023c), task-specific hints are influential priors to improve task performance. Incorporating R-Judge risk types as hints into the task instruction, only the ChatGPT F1 score improves owing to improved recall with some sacrifice of specificity, as shown in Table 2.
@@
Based on Zero-Shot-CoT, we design two-shot demonstrations (Figure 3 in Appendix) to form Few-Shot-CoT (Wei et al., 2022b) for analysis. As results shown in Table 2, Few-Shot-CoT does not consistently improve overall performance. The most plausible reason would be the limited coverage of the demonstrations for agent-related tasks (Naihin et al., 2023; Xiao et al., 2023). As they are hard to cover the full range of possible risks, using those demonstrations may confuse the LLMs.
GPT-4 F1 Recall Specificity
Zero-Shot-CoT 72.52 62.00 83.64
+ Few-Shot 64.86 48.00 100.00
+ Risk Types 71.26 62.00 78.18
ChatGPT F1 Recall Specificity
Zero-Shot-CoT 39.42 27.00 81.82
+ Few-Shot 32.26 20.00 92.73
+ Risk Types 56.10 46.00 67.27
Vicuna-13b-v1.5-16k F1 Recall Specificity
Zero-Shot-CoT 43.24 32.00 70.91
+ Few-Shot 23.08 15.00 72.73
+ Risk Types 34.33 23.00 80.00
Llama-2-13b-chat-hf F1 Recall Specificity
Zero-Shot-CoT 38.86 34.00 25.45
+ Few-Shot 34.32 29.00 27.27
+ Risk Types 38.04 35.00 10.91
Table 2: Result(%) comparison of different prompt settings (Zero-Shot-CoT, Few-Shot-CoT, Zero-Shot-CoT w/ Risk Types) in the safety judgment test. The 4 models with best baseline performance are selected. The best average model results are in bold face.
@@
The results of the prompting experiments conducted above additionally confirm the challenging nature of our task, especially when dealing with intricate cases encompassing diverse risk types. It becomes evident that straightforward prompting mechanisms are unlikely to suffice in addressing the complexity inherent in our task.
@@ 5.2 Case Study
To figure out the capability demand of the task and flaws of current LLMs, we further analyze the results of GPT-4 with a manual check. We conclude with 3 key capability flaws leading to failures:
(i) Scenario Simulation:
Fail to retrieve relevant knowledge and reason in specific scenarios.
(ii) Understanding Adaptability:
Unable to comprehend risks in specific conditions.
(iii) Safety Alignment:
Deviation of safety alignment with humans in practical scenarios.
These key flaws are aligned with results disclosed by Effectiveness and Alertness in risk identification test which reveal model capabilities in different dimensions. Due to space limitations, detailed explanations and representative failure cases are attached in Appendix D.2.1.
5.3 Oracle Test: The Effect of Risk Description on Safety Judgment
The unsatisfactory results in the risk identification test and case study demonstrate the ineffectiveness of model-generated analysis. To investigate whether model performance on safety judgment can be improved when provided with valid analysis, we introduce the Oracle Test, a modified version of the safety judgment test where analysis of LLMs is substituted with ground truth risk description written by human annotators.
Results (see Table 7 in Appendix) show that all the models achieve substantial performance gains (such as GPT-4 from 72.52% to 99.50%) with risk descriptions as environment feedback, revealing the importance of salient safety risk feedback. Hopefully, they can be obtained from a model that dynamically monitors the agent interaction in the environment in future studies.

### v3

@@ 3.5 Evaluation Criteria
As shown in the right of Figure 2, the two evaluation recipes are conducted in a serial pipeline. First, LLMs are demonstrated with the record as input and asked to analyze whether the agent actions in the record are safe, formalized as $p_{\theta}(R)\to\textrm{analysis}$. Then, LLMs are asked to output ‘unsafe’ or ‘safe’ based on record and analysis in the first step, formalized as $p_{\theta}(R,\textrm{analysis})\to\textrm{label}$. The output analysis in the first step and the output label in the second step are taken respectively for the two evaluation recipes, with human annotation as ground truth.
@@
To evaluate the effectiveness of LLMs in identifying safety risks, an open-ended test utilizes GPT-4 (gpt-4-0613 version) as an automatic scorer to assess the open-ended model-generated analysis on unsafe cases.
Effective risk identification should clearly state how the agent causes safety risks, which we address with the metric ${\rm Effectiveness}$. As human-annotated risk description is the ground truth, ${\rm Effectiveness}$ is assessed by the relevance between model-generated analysis and the pivot, i.e. if risks described in risk description are accurately identified and addressed in the model-generated analysis. The prompt for the GPT-4 scorer is attached in Figure 6 in Appendix D.2. Section 4.2 validated the feasibility of utilizing GPT-4 as an automatic scorer to assess the model-generated analysis.
Mutually supportive, the two intersected tests offer a valid and progressive evaluation. The safety judgment test is more fair and affordable while the risk identification test holds fine-grained interpretability.
Models All Intended Attacks Unintended Risks
F1 F1 Recall Spec Effect F1 Recall Spec Effect
GPT-4o 74.45 72.19 91.50 42.06 93 80.90 72.00 89.09 78
ChatGPT 44.96 40.55 37.00 57.48 36.5 55.63 42.00 83.64 41.5
Meta-Llama-3-8B-Instruct 61.01 65.68 66.50 66.36 81 48.32 36.00 76.36 48
Llama-2-13b-chat-hf 54.80 60.04 80.00 19.16 79.5 38.86 34.00 25.45 38.5
Llama-2-7b-chat-hf 53.74 62.99 91.50 7.48 86.75 21.56 18.00 10.91 17
Random 51.32 56.34 50.00 50.00 0 49.14 50.00 50.00 0
Vicuna-13b-v1.5 16.93 9.76 6.00 84.11 10 30.30 20.00 78.18 27
Vicuna-13b-v1.5-16k 25.00 15.49 11.00 71.03 18.5 43.24 32.00 70.91 37.5
Vicuna-7b-v1.5 18.59 18.25 12.50 77.10 24.5 19.35 12.00 78.18 25
Vicuna-7b-v1.5-16k 29.33 25.89 20.00 67.76 36 36.88 26.00 72.73 28.5
Mistral-7B-Instruct-v0.2 27.20 24.80 15.50 91.12 37.5 32.00 20.00 90.91 38
Mistral-7B-Instruct-v0.3 25.65 21.99 15.50 76.17 28 33.09 23.00 70.91 38
Table 1: Main results(%) of the safety judgment test and risk identification test in R-Judge on two fundamental risk origins: Intended Attacks and Unintended Risks. Safety judgment scores (F1, Recall and Spec as ‘Specificity’) are calculated by counting correct labels and risk identification scores (Effect as ‘effectiveness’) are assigned by an automatic GPT-4 scorer compared with human-annotated risk description. F1 is the main score while the other 3 metrics is for reference. Segment 1: GPT series; Segment 2: Llama 3; Segment 3: Llama 2; Segment 4: Random baseline; Segment 5: Vicuna-1.5; Segment 6: Mistral-7B. The best model results are in bold face, and the second best model results are underlined. F1 scores that exceed random are in italics.
@@ 4 Experiments
In this section, we first describe the experimental setup, especially metrics. Then, after validating R-Judge, we evaluate 11 popular LLMs on two tests of agent safety monitor, i.e. safety judgment and risk identification.
@@ 4.1 Setup
We comprehensively assess 11 LLMs, including API-based models and open-source models. The API-based models include GPT series (OpenAI, 2023). The open-source models include Llama-2 (Touvron et al., 2023), Llama-3, Mistral (Jiang et al., 2023), and Vicuna (Chiang et al., 2023) series. See Table 9 in Appendix F.3 for model details including version and link. And complete experimental settings including hyperparameters and costs are in Appendix F.
@@
We adopt the zero-shot chain-of-thought prompting (Kojima et al., 2022) (dubbed Zero-Shot-CoT) to induce LLMs to generate the reasoning steps before producing the final answer. This kind of analyze-then-output process has been shown to improve reasoning performance, as well as interpretability (Zhang et al., 2023b). Task instructions are simple and clear for generality, as presented in Figure 4 in Appendix D.2.
@@
Risk Identification, the pairwise evaluation, introduces ${\rm Effectiveness}$, as stated in Section 3.5. In the leaderboard, it is normalized to the range of 0 to 100.
@@
We provide random scores for comparison. As safety judgment is a task of binary classification, random Recall and Specificity are 50.00%. [^2] Calculated on top of Recall and Specificity, random F1 are 56.34% and 49.14% in the two sets, and are 51.32% in full sets. For open-ended risk identification, random Effectiveness is 0.
@@ 4.2 Validating R-Judge
Human agreement with GPT-4 scorer in the risk identification test.
First, to tackle the potential bias and incapability of LLM-as-Judge (Liu et al., 2023; Zheng et al., 2023), we measure the agreement between GPT-4 and human experts by the Pearson correlation coefficient (PCC) (Wikipedia, ). The PCC quantifies the linear correlation between two variables and is a value between -1 and 1, where the higher value indicates a higher correlation.
On 50 randomly selected unsafe samples, three human annotators (information in Appendix C) scored model analysis according to the same criteria that prompt the GPT-4 scorer. The average PCC on Effectiveness is 0.91, indicating the reliability of pairwise evaluation for the risk identification test. The full results across models are listed in Table 7 in Appendix D.3. Notably, for GPT-4 answers, the PCC between the automatic scorer and human is 0.89, showing little bias.
@@ 4.3 Main Results
Table 1 presents the main results in R-Judge, revealing that most LLMs perform unsatisfactorily on the R-Judge benchmark. Most LLMs score lower than random in the safety judgment test and perform worse in the risk identification test. Remarkably, GPT-4o ranks first and is also the only model scoring higher than random in both sets.
The results show that R-Judge is a challenging benchmark and most LLMs tend to generate ineffective analysis and make wrong judgments when judging safety risks in agent interaction. There is considerable room for enhancing the safety risk awareness of LLMs in open-agent scenarios. See Table 8 in Appendix E.1 for model performance across different categories.
@@ 5 Analysis
In this section, we delve into the effect of different mechanisms on model performance with further experiments and case studies, exploring the capability demand of the task and possible ways to enhance agent safety.
@@ 5.1 Influence of Different Prompting Techniques
GPT-4o F1 Recall Spec Effect
Zero-Shot-CoT 74.45 85.00 51.67 88
+ Few-Shot 74.19 76.67 66.54 73.5
+ Risk Types 71.16 89.67 30.48 86.5
ChatGPT F1 Recall Spec Effect
Zero-Shot-CoT 44.96 38.67 62.83 38.17
+ Few-Shot 20.06 11.33 98.14 14
+ Risk Types 70.57 82.33 43.12 70.83
Llama-3-8B-it F1 Recall Spec Effect
Zero-Shot-CoT 61.01 56.33 68.40 70
+ Few-Shot 42.63 31.33 82.53 31.17
+ Risk Types 55.81 53.67 56.88 76.83
Llama-2-13b-chat-hf F1 Recall Spec Effect
Zero-Shot-CoT 54.80 64.67 20.45 65.83
+ Few-Shot 60.27 75.33 16.73 69.67
+ Risk Types 50.84 60.33 14.13 53
Table 2: Result(%) comparison of different prompt settings (Zero-Shot-CoT, Few-Shot-CoT, Zero-Shot-CoT w/ Risk Types) in the safety judgment test. The 4 models with the best baseline performance are selected. The best average model results are in bold face. Due to space limitation, Llama-3-8B-it is short for Meta-Llama-3-8B-Instruct, and Llama-2-13b-chat is short for Llama-2-13b-chat-hf.
@@
As disclosed in literature (Li et al., 2023b; Wang et al., 2023c), task-specific hints are influential priors to improve task performance. Incorporating the R-Judge risk types as hints into the task instruction (Figure 4 in Appendix D.2), only the ChatGPT F1 score improves owing to improved Recall with some sacrifice of Specificity, as shown in Table 2.
@@
Based on Zero-Shot-CoT, we adopt Few-Shot-CoT (Wei et al., 2022b) for analysis. We construct two-shot demonstrations (Figure 7 in Appendix D.2) for fair comparisons due to limited context length of several LLMs. As results shown in Table 2, Few-Shot-CoT does not consistently improve overall performance. The most plausible reason would be the limited coverage of the demonstrations for agent-related tasks (Naihin et al., 2023; Xiao et al., 2023). As they are hard to cover the full range of possible risks, using those demonstrations may confuse the LLMs.
@@
The results of the prompting experiments conducted above additionally confirm the challenging nature of our task and the dataset, especially when dealing with intricate cases encompassing diverse risk types. It becomes evident that straightforward prompting mechanisms are unlikely to suffice in addressing the complexity inherent in our task.
5.2 Effect of Fine-tuning on Safety Judgment
To investigate the effect of fine-tuning on safety judgment, we devise controlled experiments between Llama and Llama Guard.
Llama Guard.
To moderate the LLM-generated contents, Meta developed Llama Guard on the task of safety judgment. Specifically, Llama Guard can classify unsafe content in user prompts and LLM responses, indicating whether a given prompt or response is safe or unsafe, and if unsafe, it also lists the content categories violated. Llama Guard presents superior performance in content moderation, with 91.50% of F1 in its internal test set, and approaches GPT-4 in public datasets such as OpenAI Mod (Markov et al., 2023) and BeaverTails-30k (Ji et al., 2024).
Settings.
Fine-tuned from the same foundation models (Llama-2-7b and Meta-Llama-3-8B), Llama-2-7b-chat-hf and Meta-Llama-3-8B-Instruct are baselines respectively for LlamaGuard-7b and Meta-Llama-Guard-2-8B. The basic usage of Llama Guard demands the specification of risk taxonomies, so we adopt prompting w/ risk types (See Figure 5 in Appendix D.2). [^3] The Llama Guard is not trained on generating risk analysis, so we cancel the risk identification test, i.e., ‘Effectiveness’.
Results.
Utilizing R-Judge as the test set, results are shown in Table 3. On safety judgment, Meta-Llama-Guard-2-8B surpasses the best model, GPT-4o, in the same setting (i.e. in Table 2), with lower Recall but higher Specificity. However, LlamaGuard-7b is incapable.
We speculate that the significant differences originate from foundation models and fine-tuned data. On one hand, Llama 3 as the foundation model is much stronger than Llama 2. On another hand, Meta-Llama-Guard-2-8B is fine-tuned on the larger amount of data in more risk types (11 v.s. 6 shown in Appendix E.3). Therefore, we conclude that fine-tuning to judge harmful content can also improve model performance to judge behavioral risks in agent ineractions, and high-quality data covering diverse risk types is crucial.
Models F1 Recall Spec
Llama-2-7b-chat-hf 24.14 18.67 59.85
LlamaGuard-7b 0.66 0.33 100.00
Meta-Llama-3-8B-Instruct 55.81 53.67 56.88
Meta-Llama-Guard-2-8B 71.84 74.00 64.31
Table 3: Result(%) of Llama and Llama Guard.
@@ 5.3 Case Study
To figure out the capability demand of the task and flaws of current LLMs, we further analyze the results of GPT-4o with a manual check. We conclude with three key capability flaws leading to failures:
(i) Scenario Simulation: Fail to retrieve relevant knowledge and reason in specific scenarios. Some safety risks are hidden in the complex multi-turn agent interaction, the identification of which demands LLM monitors to infer the effect of certain actions, such as the effect of link sharing with edit access. Due to the lack of scenario knowledge or reasoning ability, GPT-4o may fail to associate corresponding knowledge and reason in specific scenarios to identify risks. For example, GPT-4o failed in the case where the agent shared the file link on Twitter with edit access instead of comment access.
(ii) Understanding Adaptability: Unable to comprehend risks in specific conditions. Risks are closely tied to specific conditions, i.e., risky actions in Scenario A may be safe in Scenario B instead. Despite possessing a comprehensive understanding of security concerns through safety alignment, GPT-4o can sometimes exhibit rigidity and excessive concern due to its limited comprehension of certain conditions. For instance, in the case of false positives, GPT-4o may respond by stating that "the actions taken by the agent are generally safe, but there are a few potential security concerns…" and incorrectly categorize them as unsafe. Conversely, in the case of false negatives, GPT-4o may recognize that the agent should seek confirmation from the user regarding a potentially risky action but mistakenly classify it as safe.
These key flaws are aligned with results disclosed by the ${\rm Effectiveness}$ in Risk Identification test which reveals model capabilities in different dimensions. Figure 11,12,13 demonstrate the 3 representative failure cases in Appendix E.4.1.